# Retrieving snow depth distribution by downscaling ERA5 Reanalysis with ICESat-2 laser altimetry

### v1

@@ 2.4 Validation data
    The ALS survey: The survey by Melvold and Skaugen [55] provides 2 m gridded snow depths data for two winters over Hardangervidda. The survey encompassed six flight lines apart in 10 km intervals, each extending 80 km in an east-west direction with a crossline scanning width of 500 m (Figure 1). The data were collected between 3-21 April 2008, 21-24 April 2009, and 21 September 2008 (snow-free reference). During the autumn collection period, the ground was in nearly bare condition except for few perennial snow patches[55]. The snow depth maps were regridded (averaged) to 10 m spatial resolution for this study.
@@
    Sentinel-2 satellite imagery [L2A; 10] was used for visual data quality checks and validation of the presence/absence of snow in the Lake Møsvatn area (Figure 1).
Figure 2: Flow chart of the snow depth retrieval and downscaling-calibration scheme. First, snow depth is extracted by elevation differencing. This step includes co-registration and bias correction of the DEM (blue box). Consequently, a tree-structure-based regressor for downscaling is trained and implemented to predict local variability of snow depth in any location and at any time (green box). The resulting snow depth map time series are validated over time and in space (orange box). Satellite graphics source: NASA.gov
@@ 3.1 ICESat-2 snow depth retrieval
Snow depth ($SD_{IS2}$) was derived from ICESat-2 high-resolution elevation measurements through an elevation differencing workflow. ICESat-2 ATL08 data were categorized into snow-on ($IS_{snow}$) and snow-free segments based on attributes and flags present in the ATL08 data: the snow mask (segment_snowcover) from the NOAA daily snow cover product [67], as well as the (brightness_flag). Snow-free segments were used for DEM co-registration and subsequent bias correction (Figure 2, see 3.1.1, 3.1.2 ). The correction involved estimating the discrepancy ($\hat{\Delta_{h}})$ between the reference DEM ($DEM_{ref}$) and ICESat-2 snow-free measurements:
@@
Moving terrain such as water surfaces and permanent ice were excluded from the analysis using the segment_landcover reference mask from the ATL08 product [based on the Copernicus Global Land Cover dataset at 100-meter spatial resolution, 67].
The snow depth retrieved at each point (Easting, Northing, time) provides spatially and temporally incomplete information on snow dynamics. To overcome this, we resample snow depth from ERA5 Land at daily resolution and interpolated linearly for each desired point (E, N, t), differentiating with the retrieved snow depth, to create a localized dependent variable, denoted as subgrid variability ($Y$) (Figure 2). This variable represents the deviation from the aggregated mean snow depth at each point, used in the subsequent downscaling model (Section 3.2). The downscaling model predicts this subgrid variability, applying it to the ERA5 Land monthly snow depth data to generate snow depth time series. Due to the inherent nature of the downscaling model, raw outputs are biased towards average values while extreme values are underrepresented. Thus, a calibration step (Section  3.2.3) was added to better represent the full snow depth distribution. Data processing was done in Python with custom scripts that are available in a public GitHub repository (see Section  Data availability) with the help of Xarray [33] and Pandas [81]. The retrieved snow depth for Norway, spanning from October 2018 to October 2022 is available on Zenodo (DOI: 10.5281/zenodo.10048875).
@@ 3.1.1 Co-registration
Co-registration is a crucial step to align elevation datasets, with ICESat-2 snow-free data serving as a highly precise and spatially consistent reference. We used a computationally efficient gradient descent-based co-registration algorithm (for details see Supplementary Section A). The process, facilitated by the open-source xDEM tool [88], was applied to each DEM tile individually, ensuring accurate alignment across all datasets.
@@ 3.1.2 Bias correction
Bias correction estimates vertical discrepancies ($\hat{\Delta_{h}})$ between DEMs and ICESat-2 data. Given that DEMs are often patched together from multiple datasets and various sensors, captured in different seasons and at different resolutions, they must be cautiously used as a reference ground surface [35]. Magruder et al. [49] used ICESat-2 elevations to correct DEMs, taking canopy and slope into account. On the other hand, Tian and Shan [82] and Enderlin et al. [16] found that ICESat-2 ATL08 data underestimates terrain height when compared to the reference DEMs over steep terrain, and thus proposed a slope-dependent bias correction. Our study does not assert which dataset represents ground truth most accurately but focuses on quantifying the discrepancies so that we can exclude it from snow depth elevation differencing. Essentially, estimating the discrepancies ($\hat{\Delta_{h}})$ between two datasets is a regression problem that can be described with physiographic parameters, vegetation conditions, and quality metrics from ICESat-2. We employ the XGBoost [^3] The XGBoost (version 2.0.0) library can be accessed at https://xgboost.readthedocs.io/ regression model [8], a gradient-boosted decision tree (GBDT) algorithm. The regression model trained based on ICESat-2 snow-free measurements, is later used to predict elevation discrepancies ($\hat{\Delta_{h}})$ for all other DEM grid cells where no ICESat-2 snow-free data are available (Supplementary Section B). This results in bias-corrected snow depth measurements for all ICESat-2 snow-on data points. The bias correction successfully removed a significant negative skewness in the elevation difference histograms for snow-free data points (Figure S.4). In addition, considerable dependence on first and second-order DEM derivatives (slope and curvature) was detected and removed. After bias correction, NMAD values between ICESat-2 snow-free data points and DEM datasets are reduced from 0.66 m to 0.48 m (DTM1), and from 1.87 m to less than 0.62 m (FAB) (Figure S.4).
@@ 3.2 Snow depth downscaling-calibration
A second XGBoost regression model is employed to downscale ERA5 Land snow depths. XGBoost has demonstrated its effectiveness in downscaling tasks, such as total water storage anomaly from satellite gravimeter [1], precipitation [91] or wind speed [34]. We train the XGBoost regression model on (bias-corrected) ICESat-2 snow depth measurements and use a comprehensive set of topo-climatic features, including snow depth from ERA5 Land (sde_era), east, north, elevation (h_te_best_fit), slope, aspect, topographic position index (TPI, Weiss 87; see section 3.2.1), curvature, planform curvature (planc), profile curvature (profc), cumulative wind-aspect factor ($W_{uf}$, see section 3.2.2) and month of the year. These predictors offer valuable insight into the physical conditions governing snowpack dynamics. For instance, slope and curvatures are basic metrics governing snow accumulation [20]. We compute these terrain attributes using xDEM [88] based on DTM10 with Zevenbergen & Thorne algorithm [89] at 10 m resolution, consistent with the model’s output resolution. The time-varying wind fields are extracted from ERA5 Land (monthly data). We assume the stationarity of ERA5 land products, allowing us to apply the downscaling model trained on snow observations from 2018 to 2022 to other periods.
The XGBoost regression model uses decision trees in parallel structure to capture nonlinear relationships between snow depth subgrid variability and topo-climatic features. During the training, the structure and splits of the trees are guided by the goal of minimizing the prediction error on certain loss functions. Two types of loss functions are employed in separate model versions with slightly different purposes: (1) Square error (reg:squareerror) as a loss function to estimate the conditional mean of the target variable, this was the main method used to generate the spatially distributed maps, and (2) Quantile regression (reg:quantileloss) to give probabilistic predictions, such as Q50 (median), Q25 and Q75, which are used for point-based downscaling at weather station locations, to gain insights into the uncertainties of downscaling over time by validating predictions against in-situ observation. The adoption of quantile regression is inspired by its successful application in probabilistic forecasting [54, 90], where it is valued for not requiring a certain distribution of the target variable (for example a normal distribution) and its robustness to outliers. This makes it an ideal choice for our downscaling task as it is capturing the full range of possible snow depth values (difficult to capture with a prescribed distribution, many extreme values) under varying climatic conditions.
@@ 3.2.1 Topographic position index (TPI)
The TPI is a metric used to access slope position and classify various landforms. It quantifies the difference between the elevation of a central pixel and the average elevation of its neighbouring pixels (3 x 3 pixels). A TPI value of zero or near zero indicates a flat or nearly continuous slope. Positive TPI values suggest that the central pixel is significantly higher than the surrounding areas, forming a ridge or hill. Inversely, negative TPI values indicate that the central pixel is notably lower than its neighbouring areas, signifying a valley. TPI has proven effective in predicting snow distribution in alpine environments [75]. To represent landforms at different scales, we used two additional indices: tpi_9 (calculated in 9 x 9-pixel windows, equivalent to 90 m x 90 m) and tpi_27 (270 m x 270 m).
@@ 3.2.2 Cumulative wind-aspect factor
Figure 3: Quantifying the relationship between wind, aspect and snow redistribution. (a) The prevailing wind, e.g. from SE, results in a negative value on the windward side (snow erosion, $W_{uf_{negative}}$, brown) and positive values on the lee side (snow deposition, $W_{uf_{positive}}$, blue). Eight cardinal directions are plotted but the function works on any wind direction. (b) Research suggests that as wind speed increases, the transport rate increases exponentially with a power of 3 (adapted from Li et al. 41).
@@ 3.2.3 Calibration
In our study, the XGBoost downscaling model is tasked with predicting snow depth under a variety of conditions. We observed that XGBoost tends to produce a conservative estimate close to the mean, likely because the feature set does not consistently explain what critical conditions lead to extreme snow depths. This conservative tendency is a natural outcome of the model’s objective to minimize overall prediction error, often resulting in a distribution that skews towards average conditions and under-represents the extremes (known as scaling bias, further discussed in 5.2). The calibration step scales the modelled snow depth values so that their distribution matches the locally observed snow depth distribution, using quantile mapping [7, 42]. Thereby, while preserving the relative snow depth predictions of the downscaling model, the calibration removes the scaling bias for each specified quantile by a scaling factor calculated from the control points, ensuring a good representation of local snow depth distribution:
@@
Here, $x_{adj}$ denotes the calibrated prediction. For a given percentage, as determined by the cumulative distribution function (CDF) of the XGBoost downscaling model output $F_{M}(x)$ and that of the observational control points $F_{O}(x_{adj})$, the discrepancy between two distribution’s quantiles is encapsulated by the scaling factor ($\Delta$).
@@ 3.3 Model interpretation and assessment
To interpret the contribution of topo-climatic features in our tree-based models, we employed the SHAP values [48], a metric derived from game theory facilitating the understanding of the relative contribution of the model predictors (a.k.a. features). The SHAP value corresponds to the contributions of each feature to individual predictions. A feature can either play for or against the prediction (positive or negative SHAP value), and the magnitude of its SHAP value shows how significant its role is. The sum of all these contributions, plus the base prediction, gives us the final prediction.
For fair comparisons, we aggregated all high-resolution validation datasets to the same resolutions. Model validation was conducted across two different scales. At the microscale/ site scale, we captured snow distribution over typical micro-terrain features such as hills and gullies using the original model output resolution of 10 m. At the mesoscale, we aggregated data into a 100 m x 500 m grid, reflecting the 500 m width of the ALS survey swath, with 100 m intervals in the transection direction to enable effective depiction of snow variation across different aspects of hills and ridges. Successful prediction would result in a (near-) perfect data match in probability distribution and ranking correlation. To quantify downscaling performance, we employ four key statistical metrics. RMSE and $R^{2}$ scores evaluate the overall accuracy and fitness of the model. Spearman’s rank correlation coefficient ($\rho$) is used as a fidelity metric, with a high $\rho$ indicating good similarity in spatial distribution. The Kolmogorov-Smirnov D statistic (KSD) quantifies the degree of probability distribution matching, with KSD = 0 indicating a perfect match.
@@
where $z(x_{i})$ represents the snow depth at location $x_{i}$. The variogram indicates the rate at which correlation decreased with distance. By fitting variograms to the sum of the spherical model and Gaussian model from short to long ranges (as done by Rolstad et al. [76], Hugonnet et al. [35]), we identify spatial correlation of snow distribution at different scales.
@@ 4.1 Mesoscale snow depth variability
Figure 4 shows the snow depth maps for April 2008. Model input data, i.e., linearly interpolated snow depth from ERA5 Land (a), only represents large-scale variability. After downscaling, the fine-scale snow depth variability aligns with the topography both at the microscale (10 m, b) and aggregated 1 km (d). In comparison, seNorge data (c) at the same spatial resolution appears smoother with less spatial variability and overestimates snow depth in the western mountains. The differences between the two data sets are smaller for April 2009 where there was generally less snow in Southern Norway (see Supplementary Section E).
@@
Figure 4: Spatial distribution of snow depth in the Hardangervidda area, April 2008. (a) Downscaling model input, linearly interpolated snow depth from ERA5 Land. (b) Downscaled snow depth output at 10 m resolution. (c) seNorge snow depth at 1 km resolution. (d) Downscaled snow depth aggregated to 1 km resolution.
Figure 5: Snow depth profiles across Hardangervidda in April 2008 at mesoscale along six 500 m-wide ALS flight lines in W–E direction, from north to south (panels a-f). ALS/downscaled data points are aggregated to 100 m x 500 m cells.
Figure 5 shows snow depth transects at the mesoscale for April 2008. The ALS snow survey, downscaled output (based on DTM1, after calibration) and elevation are aggregated (averaged) to a resolution of 100 m x 500 m (owing the ALS survey’s 500 m transect width). All datasets follow a decreasing trend from snow depths exceeding 6 m to the West (close to the coast) to around 2 m to the East (far from the coast).
@@
The downscaled snow depths reveal impressive details in snow depth spatial variability that corresponds well to the observed snow depth magnitude and spatial variability captured by the ALS snow survey. Among the six flight lines, the calibrated downscaling model scored $R^{2}$ between 0.74 and 0.88, and RMSE values from 0.64 m to 0.43 m. The results are equally good for the 2009 ALS snow survey (Figure S.8 ), with low RMSE values of 0.37 m to 0.58 m that can be attributed to generally lower snow depths in 2009. In comparison, $R^{2}$ values of the (coarser-resolution) seNorge model data compared to the ALS survey are lower (0.27 to 0.55) and RMSE values higher (0.91 m to 1.36 m). The seNorge model data shows little spatial variability, is less sensitive to topographic features and orographic barriers, and shows about 40 % overestimation on the western mountain ridges but less to the East. Melvold and Skaugen [55] attributed the overestimation of the seNorge model to its strong reliance on weather stations located in low-elevation areas not representative for high mountains.
@@
Upon closer inspection, ALS ground truth data exhibits many spikes on east-facing slopes, likely resulting from wind redistribution and gravity processes. Our downscaling model acknowledges the importance of the aspect and wind-driven snow accumulation factors (Figure 9 c,g) but does not always reproduce corresponding snow distribution in the correct places. ERA5 land snow depth input data exhibit a north-south systematic bias, shifting from overestimation to underestimation relative to the ALS data (Figure 9 d). The model correctly removes this bias for the northernmost flight line in the eastern part of the transect, where ALS survey and downscaled output match well, but over-corrects snow depths for flight lines (b)–(d), resulting in snow depth underestimation compared to the ALS survey. Global DEMs perform similarly to those obtained using the Norwegian high-resolution datasets DTM1 and DTM10 at the mesoscale when compared to the ALS data on Hardangervidda (Figure S.7), with no visually identifiable systematic biases in snow depth maps.

### v2

@@ 2.4 Validation data
    The ALS survey: The survey by Melvold and Skaugen [60] provides 2 m gridded snow depths data for two winters over Hardangervidda. The survey encompassed six flight lines apart in 10 km intervals, each extending 80 km in an east-west direction with a crossline scanning width of 500 m (Figure 1). The data were collected between 3-21 April 2008, 21-24 April 2009, and 21 September 2008 (snow-free reference). During the autumn collection period, the ground was in nearly bare condition except for few perennial snow patches[60]. The snow depth maps were regridded (averaged) to 10 m spatial resolution for this study. Figures 6 and 5show snow depth data for April 2008 for parts of flight line b (indicated in Fig. 1).
@@
    Sentinel-2 satellite imagery [L2A; 11] was used for visual data quality checks and validation of the presence/absence of snow in the Lake Møsvatn area (Figure 1).
@@ 3.1 ICESat-2 snow depth retrieval
Snow depth ($SD_{IS2}$) was derived from ICESat-2 high-resolution elevation measurements through an elevation differencing workflow. ICESat-2 ATL08 data were categorized into snow-on ($IS_{snow}$) and snow-free segments based on attributes and flags present in the ATL08 data: the snow mask (segment_snowcover) from the National Oceanic and Atmospheric Administration (NOAA) daily snow cover product [73], as well as the (brightness_flag). Snow-free segments were used for DEM co-registration and subsequent bias correction (Figure 2, see 3.1.1, 3.1.2 ). The correction involved estimating the discrepancy ($\hat{\Delta_{h}})$ between the reference DEM ($DEM_{ref}$) and ICESat-2 snow-free measurements:
@@
Moving terrain such as water surfaces and permanent ice were excluded from the analysis using the segment_landcover reference mask from the ATL08 product [based on the Copernicus Global Land Cover dataset at 100-meter spatial resolution, 73].
Figure 2: Flow chart of the snow depth (SD) retrieval and downscaling-calibration scheme. First, snow depth is extracted by elevation differencing. This step includes co-registration and bias correction of the DEM (blue box). Consequently, a tree-structure-based regressor for downscaling is trained and implemented to predict local variability of snow depth in any location and at any time (green box). The resulting snow depth map time series are validated over time and in space (orange box). For full definitions of acronyms, see Section 3. Satellite graphics source: NASA.gov
The snow depth retrieved at each point (Easting, Northing, time) provides spatially and temporally incomplete information on snow dynamics. To overcome this, we obtained daily maximum snow depth from ERA5 Land and interpolated linearly in space for each desired point (E, N, t). The ICESat-2 snow depth is then subtracted from the interpolated ERA5 Land snow depth to create a localized dependent variable, denoted as subgrid variability ($Y$) (Figure 2). This variable represents the deviation from the aggregated mean snow depth at each point, used in the subsequent downscaling model (Section 3.2). The downscaling model predicts this subgrid variability, applying it to the ERA5 Land monthly snow depth data to generate snow depth time series. Due to the inherent nature of the downscaling model, raw outputs are biased towards average values while extreme values are underrepresented (i.e. conditional bias). Thus, a calibration step (Section  3.2.3) was added to better represent the full snow depth distribution. Data processing was done in Python with custom scripts that are available in a public GitHub repository (see Section  Data availability). The processing relies on the libraries Xarray [36] and Pandas [91]. The retrieved snow depth for Norway, spanning from October 2018 to October 2022 is available on Zenodo (DOI: 10.5281/zenodo.10048875).
@@ 3.1.1 Co-registration
Co-registration is a crucial step to align elevation datasets, with ICESat-2 snow-free data serving as a highly precise and spatially consistent reference. We used a computationally efficient gradient descent-based co-registration algorithm (for details see Supplementary Section A). The process, facilitated by the open-source xDEM tool [101], was applied to each DEM tile, ensuring accurate alignment across all datasets.
@@ 3.1.2 Bias correction
Bias correction estimates vertical discrepancies ($\hat{\Delta_{h}})$ between DEMs and ICESat-2 data. Given that DEMs are often patched together from multiple datasets and various sensors, captured in different seasons and at different resolutions, they must be cautiously used as a reference ground surface [38]. Magruder et al. [54] used ICESat-2 elevations to correct DEMs, taking canopy and slope into account. On the other hand, Tian and Shan [92] and Enderlin et al. [18] found that ICESat-2 ATL08 data underestimates terrain height when compared to the reference DEMs over steep terrain, and thus proposed a slope-dependent bias correction. Our study does not assert which dataset represents ground truth most accurately but focuses on quantifying the discrepancies so that we can exclude it from snow depth elevation differencing. Essentially, estimating the discrepancies ($\hat{\Delta_{h}})$ between two datasets is a regression problem that can be described with physiographic parameters, vegetation conditions, and quality metrics from ICESat-2. We employ the XGBoost [^3] The XGBoost (version 2.0.0) library can be accessed at https://xgboost.readthedocs.io/ regression model [9], a gradient-boosted decision tree (GBDT) algorithm. The regression model trained based on ICESat-2 snow-free measurements, is later used to predict elevation discrepancies ($\hat{\Delta_{h}})$ for all other DEM grid cells where ICESat-2 snow-free data are not available (Supplementary Section B). This results in bias-corrected snow depth measurements for all ICESat-2 snow-on data points. The bias correction successfully removed a significant negative skewness in the elevation difference histograms for snow-free data points (Figure S.4). In addition, considerable dependence on first and second-order DEM derivatives (slope and curvature) was detected and removed. After bias correction, NMAD values between ICESat-2 snow-free data points and DEM datasets are reduced from 0.66 m to 0.48 m (DTM1), and from 1.87 m to less than 0.62 m (FAB) (Figure S.4).
@@ 3.2 Snow depth downscaling-calibration
A second XGBoost regression model is employed to downscale ERA5 Land snow depths. XGBoost has demonstrated its effectiveness in downscaling tasks, such as total water storage anomaly from satellite gravimeter [1], precipitation [104] or wind speed [37]. We train the XGBoost regression model on (bias-corrected) ICESat-2 snow depth measurements and use a comprehensive set of topo-climatic features, including snow depth from ERA5 Land (sde_era), east, north, elevation (h_te_best_fit), slope, aspect, topographic position index (TPI, Weiss 100; see section 3.2.1), curvature, planform curvature (planc), profile curvature (profc), cumulative wind-aspect factor ($W_{uf}$, see section 3.2.2) and month of the year. These predictors offer valuable information into the physical driver of snowpack dynamics. For instance, slope and curvatures are basic metrics governing snow accumulation [22]. We compute these terrain attributes using xDEM [101] based on DTM10 with Zevenbergen & Thorne algorithm [102] at 10 m resolution, consistent with the model’s output resolution. The time-varying wind fields are extracted from ERA5 Land monthly data (see section 3.2.2). We assume ERA5 land products to be stationary, allowing us to apply the downscaling model trained on snow observations from 2018 to 2022 to other periods.
The XGBoost regression model uses decision trees in parallel structure to capture nonlinear relationships between snow depth subgrid variability and topo-climatic features. During training, the structure and splits of the trees are guided by the goal of minimizing the prediction error on given loss functions. Two types of loss functions are employed in separate model versions with slightly different purposes: (1) Square error (reg:squareerror) as a loss function to estimate the conditional mean of the target variable, this was the main method used to generate the spatially distributed maps, and (2) Quantile regression (reg:quantileloss) to give probabilistic predictions, such as Q50 (median), Q25 and Q75, which are used for point-based downscaling at weather station locations, to gain insights into the uncertainties of downscaling over time by validating predictions against in-situ observation. The adoption of quantile regression is inspired by its successful application in probabilistic forecasting [59, 103], for which no pre-knowledge of the target variable distribution (e.g. a normal distribution) is required while being robust to outliers. Quantile regression is therefore an ideal choice for our downscaling task as it captures the full range of possible snow depth values (e.g. extreme values) under varying climatic conditions.
@@ 3.2.1 Topographic position index (TPI)
The TPI is a metric used to assess slope position and classify various landforms. It quantifies the difference between the elevation of a central pixel and the average elevation of its neighboring pixels (3 x 3 pixels). A TPI value of zero or near zero indicates a flat or nearly continuous slope. Positive TPI values suggest that the central pixel is significantly higher than the surrounding areas, forming a ridge or hill. Inversely, negative TPI values indicate that the central pixel is notably lower than its neighboring areas, signifying a valley. TPI has proven effective in predicting snow distribution in alpine environments [84, 12]. To represent landforms at different scales, we used two additional indices: tpi_9 (calculated in 9 x 9-pixel windows, equivalent to 90 m x 90 m) and tpi_27 (270 m x 270 m).
@@ 3.2.2 Cumulative wind-aspect factor
Figure 3: Quantifying the relationship between wind, aspect and snow redistribution. (a) The prevailing wind, e.g. from SE, results in a negative value on the windward side (snow erosion, $W_{uf_{negative}}$, brown) and positive values on the lee side (snow deposition, $W_{uf_{positive}}$, blue). Eight cardinal directions are plotted but the function works on any wind direction. (b) Research suggests that as wind speed increases, the transport rate increases exponentially with a power of 3 (adapted from Li et al. 45).
@@ 3.2.3 Calibration
In our study, the XGBoost downscaling model is tasked with predicting snow depth under a variety of conditions. We observed that XGBoost tends to produce a conservative estimate close to the mean, likely because the feature set does not consistently explain what critical conditions lead to extreme snow depths. This conservative tendency is a natural outcome of the model’s objective to minimize overall prediction error, often resulting in a distribution that skews towards average conditions and under-represents the extremes (known as scaling bias, further discussed in Section 5.2). The calibration step scales the modelled snow depth values so that their distribution matches the locally observed snow depth distribution, using quantile mapping [8, 46]. Thereby, while preserving the relative snow depth predictions of the downscaling model, the calibration removes the scaling bias for each specified quantile by a scaling factor calculated from the control points, ensuring a good representation of local snow depth distribution:
@@
Here, $x_{adj}$ denotes the calibrated prediction. Determined by the cumulative distribution function (CDF) of the XGBoost downscaling model output $F_{M}(x)$ and that of the observational control points $F_{O}(x_{adj})$, the discrepancy between two distribution’s quantiles is encapsulated by the scaling factor ($\Delta$) for a given ratio.
@@ 3.3 Model interpretation and assessment
To interpret the contribution of topo-climatic features in our tree-based models, we employed the SHAP values [53], a metric derived from game theory facilitating the understanding of the relative contribution of model predictors (a.k.a. features). The SHAP value corresponds to the contributions of each feature to individual predictions. A feature can either play for or against the prediction (positive or negative SHAP value), and the magnitude of its SHAP value shows the significance of the feature role. The sum of all these contributions, plus the base prediction, provides the final prediction.
For fair comparisons, we aggregated all high-resolution validation datasets to the same resolutions. Model validation was conducted across two different scales. At the micro to site scale (respectively 100 to 10 m), we captured snow distribution over typical micro-terrain features such as hills and gullies using the original model output resolution of 10 m. At the mesoscale, we aggregated data into a 100 m x 500 m grid, reflecting the 500 m width of the ALS survey swath, with 100 m intervals in the transect direction to obtain a good characterisation of snow depth variability across different aspects of hills and ridges. Successful prediction would result in a (near-) perfect data match in probability distribution and ranking correlation. To quantify the downscaling performance, we employ four key statistical metrics. RMSE and $R2$ scores evaluate the overall accuracy and fitness of the model. $R2$ score is computed using the standard implementation of the Python library Scikit-learn:
$R2=1-\frac{\mathrm{SS_{res}}}{\mathrm{SS_{total}}}$
where $\mathrm{SS_{res}}$ represents the variation in the data that the fitted model does not explain, expressed by the sum of squared residuals between the model output and measured data, and $\mathrm{SS_{total}}$ is the total variation in the data, i.e., the sum of squared residuals with regard to the mean. The $R2$ value typically ranges from 1 (perfect fit) to 0, but can be negative if the model is evaluated on different data than used for training (as in our case) and for nonlinear models typically used in machine learning approaches (e.g. XGBoost, as used in this study). A negative $R2$ value means that the model performs worse than a constant function that always predicts the mean. As R2 is sensitive to the presence of bias, Spearman’s rank correlation coefficient ($\rho$) is used as a fidelity metric, with a high $\rho$ indicating good similarity in spatial distribution. The Kolmogorov-Smirnov D statistic (KSD) quantifies the degree of probability distribution matching, with KSD = 0 indicating a perfect match. To compare the patchy snow distribution during melt-out season with binary snow cover data from satellite imagery, we use the area under curve (AUC) and Matthews correlation coefficient (MCC). For these metrics, a value of 1 indicates a perfect match, whereas values of 0.5 (AUC) and 0 (MCC) correspond to random guessing.
@@
where $z(x_{i})$ represents the snow depth at location $x_{i}$. The variogram indicates the rate at which correlation decreased with distance. By fitting variograms to the sum of the spherical model and Gaussian model for short and long ranges, respectively [following the method of 85, 38], we identify spatial correlation of snow distribution at different scales. The variograms are computed using the xDEM tool [101].
@@ 4.1 Mesoscale snow depth variability
Figure 4 shows the snow depth maps for April 2008. Model input data, i.e., linearly interpolated snow depth from ERA5 Land (a), only represents large-scale variability. After downscaling, the fine-scale snow depth variability aligns with the topography both at the microscale (10 m, b) and aggregated to 1 km (d). In comparison, seNorge data (c) at the same spatial resolution appear smoother with less spatial variability and overestimate snow depth in the western mountains. The differences between the two data sets are smaller for April 2009 where there was generally less snow in Southern Norway (see Supplementary Section E). A comparison of April 2008 snow depths at mesoscale for flight line b is shown as a transect in Figure 5. The ALS snow survey, downscaled output (after calibration), seNorge, ERA5 Land data and elevation are aggregated (averaged) to a resolution of 100 m x 500 m, owing the ALS survey’s 500 m transect width. All datasets follow a decreasing trend from snow depths exceeding 6 m to the West (close to the coast) to around 2 m to the East (far from the coast).
@@
Figure 4: Spatial distribution of snow depth in the Hardangervidda area, April 2008. (a) Downscaling model input, linearly interpolated snow depth from ERA5 Land. (b) Downscaled snow depth output at 10 m resolution. (c) seNorge snow depth at 1 km resolution. (d) Downscaled snow depth aggregated to 1 km resolution. Map coordinates are in meters UTM 33N.
The downscaled snow depths reveal impressive details in snow depth spatial variability at the mesoscale that corresponds well to the observed snow depth magnitude and spatial variability captured by the ALS snow survey. The goodness of fit is similar for the other flight strips (Figures S.7, S.8). Across the six flight lines, the calibrated downscaling models score an $R2$ of 0.81 and an RMSE of 0.53–0.57 m (Table 1). We achieve nearly as good performance at mesoscale when using the global DEMs COP30 and FAB for ICESat-2 snow depth sample retrieval rather than the Norwegian high-resolution datasets DTM1 and DTM10 (Figure S.7). There are only minimal differences between output snow depth maps and we did not detect any systematic biases, no matter which DEM was used for snow depth sample retrieval. The results are similarly good for the 2009 ALS snow survey (Figure S.8 ), with slightly lower RMSE values of 0.50 m to 0.55 m that can be attributed to generally lower snow depths in 2009.
@@
In comparison, $R2$ values of the (coarser-resolution) seNorge model data compared to the ALS survey are lower (0.49 and 0.65) and RMSE values higher (1.12 m and 0.60 m). The seNorge model shows little spatial variability, is less sensitive to topographic features and orographic barriers, and shows about 40 % overestimation on the western mountain ridges but matches ALS data in the East. Melvold and Skaugen [60] attributed the overestimation of the seNorge model to its strong reliance on weather stations located in low-elevation areas not representative for high mountains.
@@
Upon closer inspection, ALS ground truth data exhibit many spikes on east-facing slopes (Figure 5), likely resulting from wind redistribution and gravity processes. Our downscaling model acknowledges the importance of the aspect and wind-driven snow accumulation factors (see also Figure 9 c,g) but does not always reproduce corresponding snow distribution in the correct places. ERA5 land snow depth input data exhibit a north-south systematic bias, shifting from overestimation to underestimation relative to the ALS data (Figure 9 d). The model over-corrects snow depths in the Eastern part of flight line b, resulting in snow depth underestimation compared to the ALS survey.
Figure 5: Snow depth profile of flight line b across Hardangervidda as shown in Figure 1). Data is shown for April 2008 at mesoscale, i.e. averaged to 100 m x 500 m (flight strip width) cells along the flight strip length of 80 km. Model outputs of the different DEMs are nearly identical. The map strips illustrate snow depth from ERA5 Land, seNorge, ALS snow survey and downscaled model output using the DTM1 for the marked section of the datasets, in their native resolution before aggregation.
Table 1: Statistical comparison of snow depth estimates from different methods (downscaled output with snow retrievals based on different DEMs, seNorge, and ERA5 Land) against ALS snow survey data for all six flight lines (a–f) combined, for the years 2008 and 2009, respectively. $R2$, KSD and Spearmans’ $\rho$ are unit-less, RMSE is in m.
April 2008 April 2009
Dataset $R2$ KSD $\rho$ RMSE $R2$ KSD $\rho$ RMSE
DTM1 0.81 0.09 0.88 0.53 0.78 0.07 0.87 0.50
DTM10 0.82 0.12 0.88 0.56 0.79 0.08 0.88 0.513
COP30 0.80 0.10 0.88 0.57 0.75 0.07 0.84 0.554
FAB 0.81 0.11 0.89 0.56 0.77 0.07 0.86 0.54
SeNorge 0.49 0.21 0.84 1.12 0.65 0.13 0.86 0.60
ERA5 Land -0.79 0.27 0.71 0.82 -0.34 0.24 0.72 0.68